# Maker-Breaker domination game critical graphs

### v1

@@ Abstract
The Maker-Breaker domination game (MBD game) is a two-player game played on a graph $G$ by Dominator and Staller. They alternately select unplayed vertices of $G$. The goal of Dominator is to form a dominating set with the set of vertices selected by him while that of Staller is to prevent this from happening. In this paper MBD game critical graphs are introduced. Their existence is established and critical graphs are characterized for most of the cases in which the first player can win the game in one or two moves.
@@ 1 Introduction
Erdős and Selfridge introduced the Maker-Breaker game in [10]. This is a two-person game played on an arbitrary hypergraph $\mathcal{H}$. The players named Maker and Breaker alternately select an unplayed vertex of $\mathcal{H}$ during the game. Maker aims to occupy all the vertices of some hyperedge, on the other hand, Breaker’s goal is to prevent Maker from doing it. The game has been extensively researched, both in general and in specific cases, cf. the book [16], the recent paper [18], and references therein.
@@
It is known from [7, Lemma 2.3] that all the four graph invariants associated with the MBD game are monotonic for adding and deleting an edge. This motivates us to introduce MBD game critical graphs. For the classical domination game, this aspect has already been studied in [5, 17, 19, 20]. In the preliminaries, additional definitions are listed and known results about the MBD game needed later on are stated. MBD game critical graphs are formally defined in Section 3. In the same section the existence of the MBD game critical graphs is established for all four related invariants, in three of the four cases explicit constructions are provided. In Section 4 critical graphs are characterized for most of the cases in which Dominator wins the game in one or two moves. Parallel results for graphs in which Staller wins the game in one or two moves are derived in Section 5.
@@ 2 Preliminaries
Let $G=(V(G),E(G))$ be a graph. The order of $G$ is denoted by $n(G)$. For a vertex $v\in V(G)$, its open neighbourhood is denoted by $N(v)$ and its closed neighbourhood by $N[v]$. The degree of $v$ is $\deg(v)=|N(v)|$. The minimum and the maximum degree of $G$ are respectively denoted by $\delta(G)$ and $\Delta(G)$. An isolated vertex is a vertex of degree 00, a leaf is a vertex of degree $1$. A support vertex is a vertex adjacent to a leaf and a strong support vertex is a vertex that is adjacent to at least two leaves. A set $S\subseteq V(G)$ is $x$-free for $x\in V(G)$ if $x\notin S$.
A dominating set of $G$ is a set $D\subseteq V(G)$ such that each vertex from $V(G)\setminus D$ has a neighbour in $D$. The domination number $\gamma(G)$ of $G$ is the minimum cardinality of a dominating set of $G$. If $X$ is a dominating set of $G$ with $|X|=\gamma(G)$, then $X$ is a $\gamma$-set of $G$. A vertex of degree $n(G)-1$ is a dominating vertex. If $G$ is a connected bipartite graph, then its vertex $x$ is a bipartite dominating vertex if $x$ is adjacent to all the vertices of the bipartition set of $G$ which does not contain $x$. An edge of a graph $G$ is a dominating edge if it is adjacent to all the other edges of $G$.
The outcome $o(G)$ of the MBD game played on $G$ can be one of $\mathcal{D}$, $\mathcal{S}$, and $\mathcal{N}$, where $o(G)=\mathcal{D}$, if Dominator has a winning strategy no matter who starts the game; $o(G)=\mathcal{S}$, if Staller has a winning strategy no matter who starts the game; and $o(G)=\mathcal{N}$, if the first player has a winning strategy. See [8] that the fourth possible option for the outcome never happens.
@@ 3 MBD game critical graphs
In this section, we introduce MBD game critical graphs. It is known from [3] that the outcome of the MBD game of a graph $G$ may change when an edge is removed or added. It was observed in [7] that the MBD number of a graph $G$ never decreases by removing an edge and that its SMBD number never decreases by adding an edge. Hence the MBD numbers and the SMBD numbers of $G$ are monotonic with respect to the deletion and addition of an edge. This motivates us to define the MBD game critical graphs as follows.
@@
A graph $G$ is
  • $\gamma_{\rm MB}$-critical, if $\gamma_{\rm MB}(G)<\infty$ and $\gamma_{\rm MB}(G)\neq\gamma_{\rm MB}(G-e)$, for any $e\in E(G)$;
  • $\gamma_{\rm MB}^{\prime}$-critical, if $\gamma_{\rm MB}^{\prime}(G)<\infty$ and $\gamma_{\rm MB}^{\prime}(G)\neq\gamma_{\rm MB}^{\prime}(G-e)$, for any $e\in E(G)$;
  • $\gamma_{\rm SMB}$-critical, if $\gamma_{\rm SMB}(G)<\infty$ and $\gamma_{\rm SMB}(G)\neq\gamma_{\rm SMB}(G-e)$, for any $e\in E(G)$;
  • $\gamma_{\rm SMB}^{\prime}$-critical, if $\gamma_{\rm SMB}^{\prime}(G)<\infty$ and $\gamma_{\rm SMB}^{\prime}(G)\neq\gamma_{\rm SMB}^{\prime}(G-e)$, for any $e\in E(G)$.
Note that if $G$ is a graph with $\gamma_{\rm MB}(G)=\infty$, then by Lemma 2.3(i) we have $\gamma_{\rm MB}(G-e)\geq\gamma_{\rm MB}(G)=\infty$, thus it is not possible that $\gamma_{\rm MB}(G)\neq\gamma_{\rm MB}(G-e)$ holds for any edge $e\in E(G)$. Therefore we could omit the condition $\gamma_{\rm MB}(G)<\infty$ from the definition of $\gamma_{\rm MB}$-critical graphs. By the same reasoning, we could also omit the condition $\gamma_{\rm MB}^{\prime}(G)<\infty$ from the definition of $\gamma_{\rm MB}^{\prime}$-critical graphs.
Lemma 3.2
If $G$ is a graph and $e\in E(G)$, then the following properties hold.
  • (i)
    If $G$ is $\gamma_{\rm MB}$-critical, then $\gamma_{\rm MB}(G)<\gamma_{\rm MB}(G-e)$.
  • (ii)
    If $G$ is $\gamma_{\rm MB}^{\prime}$-critical, then $\gamma_{\rm MB}^{\prime}(G)<\gamma_{\rm MB}^{\prime}(G-e)$.
  • (iii)
    If $G$ is $\gamma_{\rm SMB}$-critical, then $\gamma_{\rm SMB}(G)>\gamma_{\rm SMB}(G-e)$.
  • (iv)
    If $G$ is $\gamma_{\rm SMB}^{\prime}$-critical, then $\gamma_{\rm SMB}^{\prime}(G)>\gamma_{\rm SMB}^{\prime}(G-e)$.
Proof. Assume that $G$ is $\gamma_{\rm MB}$-critical. Then $\gamma_{\rm MB}(G)\neq\gamma_{\rm MB}(G-f)$ holds for any $f\in E(G)$, hence (i) follows by Lemma 2.3(i). Statements (ii), (iii), and (iv) follow by the same argument by respectively applying Lemma 2.3(ii), (iii), and (iv). $\square$
If $G$ is $\gamma_{\rm MB}$-critical and $\gamma_{\rm MB}(G)=k$, then we say that $G$ is a $k$-$\gamma_{\rm MB}$-critical. We analogously define $k$-$\gamma_{\rm MB}^{\prime}$-critical, $k$-$\gamma_{\rm SMB}$-critical, and $k$-$\gamma_{\rm SMB}^{\prime}$-critical graphs.
It is clear that the disjoint union of $k$ copies of $K_{2}$ is a $k$-$\gamma_{\rm MB}$-critical graph. To show that there exist connected such graphs, consider graphs $G_{k}$, $k\geq 1$, constructed as follows. First, take the disjoint union of $k$ copies of $K_{2,2}$ with respective bipartitions $\{x_{i},x_{i}^{\prime}\}$, $\{y_{i},y_{i}^{\prime}\}$, where $i\in[k]$. Then add vertex $w$ and make it adjacent to $x_{i}$ and $x_{i}^{\prime}$ for $i\in[k]$. Finally, add a vertex $w^{\prime}$ and the edge $ww^{\prime}$. See Fig. 1.
@@
Figure 1: The graph $G_{k}$ which is a connected $(k+1)$-$\gamma_{\rm MB}$ critical graph
@@
We next construct (connected) examples of $k$-$\gamma_{\rm SMB}^{\prime}$-critical graphs. To this end, we recall from [3, Theorem 5.2] that for an odd $n$,
$\gamma_{\rm SMB}^{\prime}(P_{n})=\lfloor\log_{2}(n)\rfloor+1\,.$ (1)
@@
If $k\geq 1$, then $P_{2^{k}+1}$ is a $(k+1)$-$\gamma_{\rm SMB}^{\prime}$-critical graph.
Proof. Let $n=2^{k}+1$ for some positive integer $k$. Then by (1), $\gamma_{\rm SMB}^{\prime}(P_{n})=k+1$. Let $e$ be an arbitrary edge of the path. Then at least one component of $P_{n}-e$ is a path $P$ of odd order strictly smaller than $2^{k}$. Then the strategy of Staller is to select vertices from $P$. Since the order of $P$ is strictly less than $2^{k}$ it follows from [3] that $\gamma_{\rm SMB}^{\prime}(P_{n}-e)=\gamma_{\rm SMB}^{\prime}(P)<k+1$. $\square$
From Proposition 3.5 we can deduce that the disjoint union of $P_{2^{k}+1}$ and an isolated vertex is a $(k+1)$-$\gamma_{\rm SMB}$-critical graph. In the following proposition, we give an existential proof of the existence of $k$-$\gamma_{\rm SMB}$-critical graphs.
Proposition 3.6
For any integer $k>1$, there exists a $k$-$\gamma_{\rm SMB}$-critical graph.
Proof. Let $G$ be an arbitrary graph with $\gamma_{\rm SMB}(G)=k$. (It is known from [3, Theorem 3.2] that such graphs exist.) If $G$ is $k$-$\gamma_{\rm SMB}$-critical we are done. Otherwise $G$ contains an edge $e$ such that $\gamma_{\rm SMB}(G-e)=k$. Continuing the process of removing such edges from $G$ we arrive at a graph $G^{\prime}$ with $\gamma_{\rm SMB}(G^{\prime})=k$ and such that $\gamma_{\rm SMB}(G^{\prime}-e)\neq k$ for every $e\in E(G^{\prime})$. Hence $G^{\prime}$ is $k$-$\gamma_{\rm SMB}$-critical. $\square$
Note that the proof of Proposition 3.6 reveals that any graph $G$ contains a spanning subgraph $G^{\prime}$ which is is $\gamma_{\rm SMB}$-critical.
4 MBD game critical graphs with small MBD numbers
Theorem 4.1
@@ 4 MBD game critical graphs with small MBD numbers
Proof. First, assume that $G$ is connected and $1$-$\gamma_{\rm MB}$-critical. Since $G$ is $1$-$\gamma_{\rm MB}$-critical, we have $\gamma_{\rm MB}(G)=1$ and $\gamma_{\rm MB}(G-e)>1$ for every $e\in E(G)$. Since $\gamma_{\rm MB}(G)=1$, $G$ contains at least one dominating vertex. Suppose first that $G$ contains two dominating vertices $x,y$. Then for any edge $e\in E(G)\setminus\{xy\}$ it holds that $\gamma_{\rm MB}(G-e)=1$. Since $\gamma_{\rm MB}(G-f)>1$ holds for any edge $f\in E(G)$, this implies that $E(G)=\{xy\}$ and thus $G$ is isomorphic to $K_{2}$ or equivalently to $K_{1,1}$. Suppose now that $G$ has exactly one dominating vertex, say $u$. If possible, suppose that there exists an edge $e$ of $G$ that is not incident with $u$. Then $\gamma_{\rm MB}(G-e)=1$, which leads to a contradiction. Hence every edge of $G$ is incident with $u$. Thus $G$ is isomorphic to $K_{1,n},n\geq 1$.
@@
Conversely, assume that $G=K^{\prime}_{2,n}$ for some $n\geq 1$. Then $G$ has two dominating vertices and hence $\gamma_{\rm MB}^{\prime}(G)=1$. Moreover, $G-e$ has at most one dominating vertex for every $e\in E(G)$. Thus in the S-game played on $G-e$, Dominator needs at least two moves because Staller can choose the dominating vertex of $G-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}e}$ in her first move if there is such a vertex. Therefore $G$ is $1$-$\gamma_{\rm MB}^{\prime}$-critical. $\square$
@@
A connected graph $G$ is $2$-$\gamma_{\rm MB}$-critical if and only if $G$ is obtained from a star $K_{1,n}$, $n\geq 1$, with center $u$ and a $K_{2,m}$, $m\geq 2$, whose bipartition is $\{x_{1},x_{2}\}$, $\{y_{1},y_{2},\ldots,y_{m}\}$, by adding the edges $ux_{1}$ and $ux_{2}$.
@@
Figure 3: An example of connected $2$-$\gamma_{\rm MB}$-critical graph
@@
From all these cases we conclude that $G$ is 2-$\gamma_{\rm MB}$ critical. $\square$
@@
Proof. (i) Let $G$ be a connected $2$-$\gamma_{\rm MB}^{\prime}$-critical graph. Since $\gamma_{\rm MB}^{\prime}(G)=2$, $G$ has at least four vertices (note that Staller has two moves and Dominator has two moves). Hence we need to exclude all connected graphs of order $4$. Clearly, $K_{4}$ and $K_{4}-e$ are connected graphs having four vertices. But it is clear that $\gamma_{\rm MB}^{\prime}(K_{4})=\gamma_{\rm MB}^{\prime}(K_{4}-e)=1$. The remaining connected graphs with four vertices are $P_{4}$, $C_{4}$ and $K_{3}$ with a pendant edge. We can easily verify that $\gamma_{\rm MB}^{\prime}(P_{4})=\gamma_{\rm MB}^{\prime}(C_{4})=2$. Since $P_{4}$ is isomorphic to $C_{4}-e$, $C_{4}$ is not a critical graph. Let $v_{1},v_{2},v_{3},v_{4}$ be the vertices of $P_{4}$ and let $e$ be the edge $v_{2}v_{3}$. Clearly, $P_{4}-e$ is the disjoint union of two $K_{2}$’s. Therefore $\gamma_{\rm MB}^{\prime}(P_{4}-e)=2$. Hence $P_{4}$ is not a critical graph. Finally, let $H$ be a graph with vertices $x_{1},x_{2},x_{3}$ that induce a $K_{3}$ and $y$ that is adjacent to $x_{1}$, and there are no other vertices or edges in $H$. Then $\gamma_{\rm MB}^{\prime}(H-x_{1}x_{3})=2$, as $H-x_{1}x_{3}$ is isomorphic to $P_{4}$. Hence $H$ is not 2-$\gamma_{\rm MB}^{\prime}$-critical. Thus there is no connected $2$-$\gamma_{\rm MB}^{\prime}$ critical graph having four vertices.
@@
Now consider an S-game on $G$. Assume that Staller selects $v$ as her first optimal move. Therefore Dominator must play $u$ as his first reply, otherwise, Staller will win this game by selecting $u$. Since $\gamma_{\rm MB}^{\prime}(G)=2$, Dominator can dominate all the vertices of $G$ other than $u$ and $v$ in his next move. Hence there exist $a,b\in V(G)\setminus\{u,v\}$ that dominate all the vertices of $V(G)\setminus\{u,v\}$ or, equivalently, the subgraph of $G$ induced by $V(G)\setminus\{u,v\}$ has two dominating vertices. Let $e$ be an arbitrary edge incident with $v$ but not with $u$. Then $\gamma_{\rm MB}^{\prime}(G-e)=2$, a contradiction with $G$ being 2-$\gamma_{\rm MB}^{\prime}$-critical. Thus we conclude that $\delta(G)\geq 2$.
(iii) Let $G$ be connected and $2$-$\gamma_{\rm MB}^{\prime}$-critical. Since $\gamma_{\rm MB}^{\prime}(G)\neq 1$, $G$ has at most one dominating vertex. For the purpose of contradiction assume that $G$ contains a dominating vertex $u$, i.e. $\deg(u)=n(G)-1$. It is known from [6, Proposition 4.2] that $\gamma_{\rm MB}^{\prime}(G)=2$ implies that for any vertex $u$ there exists a vertex $v$ such that $\{v,v_{1}\}$ and $\{v,v_{2}\}$ are two $u$-free $\gamma$-sets. Since $\deg(v)<n-1$, there is a vertex $w$ in $G$ which is not adjacent to $v$. We consider two cases for the remaining part.
@@
In this case, $v_{1}v_{2}$ must be an edge of $G$. Let $e=uv_{1}$. Consider an S-game on $G-e$. If Staller selects a vertex from the set $\{v_{1},v_{2}\}$, then Dominator in his first move selects $u$. Then $v_{1}$ is the only still not dominated vertex. Since $v_{1}$ is dominated by both $v_{1}$ and $v_{2}$, Dominator can select one of them (depending on Staller’s move) and win in two moves. Therefore $\gamma_{\rm MB}^{\prime}(G-e){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}=}2$, a contradiction.
Now, if Staller selects $u$, then Dominator can select all the vertices from either $\{v,v_{1}\}$ or $\{v,v_{2}\}$ in his first two moves and finishes the game. Finally, if Staller selects one of the vertices in $\{v_{1},v_{2}\}$, then Dominator selects the other one. Now Dominator selects one vertex from $\{u,v\}$ (depending on the Staller’s move) and wins the game in his second move. This leads to a contradiction to the assumption that $G$ is $2$-$\gamma_{\rm MB}^{\prime}$-critical.
@@
To conclude the section we identify a large class of connected graphs which are $2$-$\gamma_{\rm MB}^{\prime}$-critical.
Theorem 4.5
@@
5 MBD game critical graphs with small SMBD numbers
In this section, we characterise MBD game critical graphs with small SMBD numbers. It is known that $\gamma_{\rm SMB}(G)=1$ if and only if $G$ has at least two isolated vertices. Let $G$ be a graph with at least two isolated vertices. It is clear that $G-e$ has again at least two isolated vertices for any edge $e$ of $G$. This indeed proves that $G$ is not $1$-$\gamma_{\rm SMB}$-critical. Hence there do not exist a $1$-$\gamma_{\rm SMB}$-critical graphs.

### v2

@@ Abstract
The Maker-Breaker domination game (MBD game) is a two-player game played on a graph $G$ by Dominator and Staller. They alternately select unplayed vertices of $G$. The goal of Dominator is to form a dominating set with the set of vertices selected by him while that of Staller is to prevent this from happening. In this paper MBD game critical graphs are studied. Their existence is established and critical graphs are characterized for most of the cases in which the first player can win the game in one or two moves.
@@ 1 Introduction
It is usually said that the Maker-Breaker game was introduced by Erdős and Selfridge in [10], but it appeared already earlier in a paper of Hales and Jewett [16]. Anyhow, this is a two-person game played on an arbitrary hypergraph $\mathcal{H}$. The players named Maker and Breaker alternately select an unplayed vertex of $\mathcal{H}$ during the game. Maker aims to occupy all the vertices of some hyperedge, on the other hand, Breaker’s goal is to prevent Maker from doing it. The game has been extensively researched, both in general and in specific cases, cf. the book [17], the recent paper [19], and references therein.
@@
It is known from [7, Lemma 2.3] that all the four graph invariants associated with the MBD game are monotonic for adding and deleting an edge. This motivates us to introduce MBD game critical graphs as respectively the graphs for which $\gamma_{\rm MB}$, $\gamma_{\rm MB}^{\prime}$, $\gamma_{\rm SMB}$, $\gamma_{\rm SMB}^{\prime}$ changes as soon as an arbitrary edge is removed/added. For the classical domination game, this aspect has already been studied in [5, 18, 20, 21]. In the preliminaries, additional definitions are listed and known results about the MBD game needed later on are stated. In Section 3, $\gamma_{\rm MB}$- and $\gamma_{\rm MB}^{\prime}$-critical graphs are formally defined. In the same section the existence of the $\gamma_{\rm MB}$- and $\gamma_{\rm MB}^{\prime}$-critical graphs is established and explicit constructions are provided for both cases. In the subsequent section 1-$\gamma_{\rm MB}$-critical, 1-$\gamma_{\rm MB}^{\prime}$-critical, and 2-$\gamma_{\rm MB}$-critical graphs are characterized, while 2-$\gamma_{\rm MB}^{\prime}$-critical graphs are characterized among bipartite graphs. Finally, $\gamma_{\rm SMB}$- and $\gamma_{\rm SMB}^{\prime}$-critical graphs are formally defined and characterized for graphs $G$ with $\gamma_{\rm SMB}(G),\gamma_{\rm SMB}^{\prime}(G)\in\{1,2\}$ in Section 5.
@@ 2 Preliminaries
Let $G=(V(G),E(G))$ be a graph. The order of $G$ is denoted by $n(G)$. For a vertex $v\in V(G)$, its open neighbourhood is denoted by $N(v)$ and its closed neighbourhood by $N[v]$. The degree of $v$ is $\deg(v)=|N(v)|$. The minimum and the maximum degree of $G$ are respectively denoted by $\delta(G)$ and $\Delta(G)$. An isolated vertex is a vertex of degree 00, a leaf is a vertex of degree $1$. A support vertex is a vertex adjacent to a leaf and a strong support vertex is a vertex that is adjacent to at least two leaves. A set $S\subseteq V(G)$ is $x$-free for $x\in V(G)$ if $x\notin S$. As usual, $K_{n}$ and $P_{n}$ denote the complete graph and the path of order $n$, while $K_{m,n}$ is the complete bipartite graph of order $m+n$.
A dominating set of $G$ is a set $D\subseteq V(G)$ such that each vertex from $V(G)\setminus D$ has a neighbour in $D$. The domination number $\gamma(G)$ of $G$ is the minimum cardinality of a dominating set of $G$. If $X$ is a dominating set of $G$ with $|X|=\gamma(G)$, then $X$ is a $\gamma$-set of $G$. A vertex of degree $n(G)-1$ is a dominating vertex. If $G$ is a connected bipartite graph, then a vertex $x$ is a bipartite dominating vertex if $x$ is adjacent to all the vertices of the bipartition set of $G$ which does not contain $x$. An edge of a graph $G$ is a dominating edge if it is adjacent to all the other edges of $G$.
The outcome $o(G)$ of the MBD game played on $G$ can be one of $\mathcal{D}$, $\mathcal{S}$, and $\mathcal{N}$, where $o(G)=\mathcal{D}$, if Dominator has a winning strategy no matter who starts the game; $o(G)=\mathcal{S}$, if Staller has a winning strategy no matter who starts the game; and $o(G)=\mathcal{N}$, if the first player has a winning strategy. See [8] that the fourth possible option for the outcome never happens. We also add that in an optimal strategy of Dominator to achieve $\gamma_{\rm MB}(G)$ or $\gamma_{\rm MB}^{\prime}(G)$, it is never an advantage for him to skip a move. Moreover, if Staller skips a move it can never disadvantage Dominator [15, Lemma 2.3]. The same holds for the games in which Staller wins [3, Lemma 2.6].
@@ 3 MBD game critical graphs
In this section, we introduce MBD game critical graphs. It is known from [3] that the outcome of the MBD game of a graph $G$ may change when an edge is removed or added. By Lemma 2.3, the MBD number of a graph $G$ never decreases by removing an edge. This motivates us to define the MBD game critical graphs as follows.
@@
If $G$ is a graph and $\tau\in\{\gamma_{\rm MB},\gamma_{\rm MB}^{\prime}\}$, then $G$ is $\tau$-critical, if $\tau(G)<\tau(G-e)$, for any $e\in E(G)$.
Note that in view of Lemma 2.3 (i) and (ii), if $G$ is $\tau$-critical, where $\tau\in\{\gamma_{\rm MB},\gamma_{\rm MB}^{\prime}\}$, then $\tau(G)<\infty$. If $G$ is $\tau$-critical and $\tau(G)=k$, where $\tau\in\{\gamma_{\rm MB},\gamma_{\rm MB}^{\prime}\}$, then we say that $G$ is a $k$-$\tau$-critical.
Let $\gamma_{\rm MB}(G)=k$. Iteratively removing edges $e$ with the property $\gamma_{\rm MB}(G-e)=k$ we arrive at a $k$-$\gamma_{\rm MB}$-critical spanning subgraph of $G$. Analogous conclusion holds for a $k$-$\gamma_{\rm MB}^{\prime}$-critical graph. So such critical graphs clearly exists. In the rest of the section we prove that also connected critical graphs exist.
Consider graphs $G_{k}$, $k\geq 1$, constructed as follows. First, take the disjoint union of $k$ copies of $K_{2,2}$ with respective bipartitions $\{x_{i},x_{i}^{\prime}\}$, $\{y_{i},y_{i}^{\prime}\}$, where $i\in[k]$. Then add vertex $w$ and make it adjacent to $x_{i}$ and $x_{i}^{\prime}$ for $i\in[k]$. Finally, add a vertex $w^{\prime}$ and the edge $ww^{\prime}$. See Fig. 1.
@@
Figure 1: The graph $G_{k}$ which is a connected $(k+1)$-$\gamma_{\rm MB}$-critical graph
@@
4 Critical graphs with small MBD numbers
@@
Proof. First, assume that $G$ is connected and $1$-$\gamma_{\rm MB}$-critical. Then $G$ contains a dominating vertex, say $u$. If possible, suppose that there exists an edge $e$ of $G$ that is not incident with $u$. Then $\gamma_{\rm MB}(G-e)=1$, which leads to a contradiction. Hence every edge of $G$ is incident with $u$. Thus $G$ is isomorphic to $K_{1,n},n\geq 1$.
@@
Conversely, assume that $G=K^{\prime}_{2,n}$ for some $n\geq 1$. Then $G$ has two dominating vertices and hence $\gamma_{\rm MB}^{\prime}(G)=1$. Moreover, $G-e$ has at most one dominating vertex for every $e\in E(G)$. Thus in the S-game played on $G-e$, Dominator needs at least two moves because Staller can choose the dominating vertex of $G-e$ in her first move if there is such a vertex. Therefore $G$ is $1$-$\gamma_{\rm MB}^{\prime}$-critical. $\square$
@@
A connected graph $G$ is $2$-$\gamma_{\rm MB}$-critical if and only if $G$ is obtained from a star $K_{1,n}$, $n\geq 1$, with center $u$ and a $K_{2,m}$, $m\geq 2$, whose bipartition is $\{x_{1},x_{2}\}$, $\{y_{1},y_{2},\ldots,y_{m}\}$, by adding the edges $ux_{1}$ and $ux_{2}$ (see Fig. 3).
@@
Figure 3: Connected $2$-$\gamma_{\rm MB}$-critical graphs
@@
From all these cases we conclude that $G$ is 2-$\gamma_{\rm MB}$-critical. $\square$
@@
Proof. (i) Let $G$ be a connected $2$-$\gamma_{\rm MB}^{\prime}$-critical graph. Since $\gamma_{\rm MB}^{\prime}(G)=2$, $G$ has at least four vertices (note that Staller has two moves and Dominator has two moves). Suppose that $G$ is 2-$\gamma_{\rm MB}^{\prime}$-critical connected graph with $n(G)=4$. Then $G$ contains $2K_{2}$ as strict spanning subgraph and thus $G$ cannot be 2-$\gamma_{\rm MB}^{\prime}$-critical.
@@
Now consider an S-game on $G$. The vertex $v$ is an optimal first move of Staller, cf. [8]. Therefore Dominator must play $u$ as his first reply, otherwise, Staller will win this game by selecting $u$. Since $\gamma_{\rm MB}^{\prime}(G)=2$, Dominator can dominate all the vertices of $G$ other than $u$ and $v$ in his next move. Hence there exist $a,b\in V(G)\setminus\{u,v\}$ that dominate all the vertices of $V(G)\setminus\{u,v\}$ or, equivalently, the subgraph of $G$ induced by $V(G)\setminus\{u,v\}$ has two dominating vertices. Let $e$ be an arbitrary edge incident with $v$ but not with $u$. Then $\gamma_{\rm MB}^{\prime}(G-e)=2$, a contradiction with $G$ being 2-$\gamma_{\rm MB}^{\prime}$-critical. Thus we conclude that $\delta(G)\geq 2$.
(iii) Let $G$ be connected and $2$-$\gamma_{\rm MB}^{\prime}$-critical. Since $\gamma_{\rm MB}^{\prime}(G)\neq 1$, $G$ has at most one dominating vertex. For the purpose of contradiction assume that $G$ contains a dominating vertex $u$, i.e. $\deg(u)=n(G)-1$. It is known from [6, Proposition 4.2] that $\gamma_{\rm MB}^{\prime}(G)=2$ implies that there exists a vertex $v\neq u$ such that $\{v,v_{1}\}$ and $\{v,v_{2}\}$ are two $u$-free $\gamma$-sets. Since $\deg(v)<n-1$, there is a vertex $w$ in $G$ which is not adjacent to $v$. We consider two cases for the remaining part.
@@
In this case, $v_{1}v_{2}$ must be an edge of $G$. Let $e=uv_{1}$. Consider an S-game on $G-e$. If Staller selects a vertex from the set $\{v_{1},v_{2}\}$, then Dominator selects the other one. Now Dominator selects one vertex from $\{u,v\}$ (depending on the Staller’s move) and wins the game in his second move. This leads to a contradiction to the assumption that $G$ is $2$-$\gamma_{\rm MB}^{\prime}$-critical. Now, if Staller selects $u$, then Dominator can select all the vertices from either $\{v,v_{1}\}$ or $\{v,v_{2}\}$ in his first two moves and finishes the game in two moves, a contradiction.
@@
To conclude the section we characterize connected bipartite graphs that are $2$-$\gamma_{\rm MB}^{\prime}$-critical.The result will follow from the following two lemmas.
Lemma 4.5
@@
Lemma 4.6
Let $G$ be a connected bipartite graph with bipartition $V_{1},V_{2}$. If $G$ is $2$-$\gamma_{\rm MB}^{\prime}$-critical, then $G$ is either $K_{2,m}$ for $m\geq 3$, or $|V_{i}|\geq 3$ for $i\in[2]$ and $G$ has exactly two bipartite dominating vertices in each $V_{i}$, $i\in[2]$ and all other vertices of $G$ are of degree $2$.
Proof. Let $V_{1}=\{v_{1,1},\ldots,v_{1,m}\}$ and $V_{2}=\{v_{2,1},\ldots,v_{2,n}\}$. First assume that $G$ is $2$-$\gamma_{\rm MB}^{\prime}$-critical. If $|V_{i}|=1$ for some $i\in[2]$, then Proposition 4.4 implies that $|V_{j}|\geq 4$ for $j\in[2]\setminus\{i\}$ and thus $G=K_{1,\ell}$, $\ell\geq 4$. Hence $\gamma_{\rm MB}^{\prime}(G)=\infty$, a contradiction with $G$ being $2$-$\gamma_{\rm MB}^{\prime}$-critical. Therefore, since $n(G)\geq 5$ by Proposition 4.4, we may without loss of generality assume that $|V_{1}|\geq 2$ and $|V_{2}|\geq 3$.
First assume that $|V_{i}|=2$ for some $i\in[2]$ and consequently $|V_{j}|\geq 3$ for $j\in[2]\setminus\{i\}$. Since $\delta(G)\geq 2$ by Proposition 4.4, any vertex $x\in V_{j}$ is adjacent to both vertices of $V_{i}$. Hence $G$ is isomorphic to $K_{2,m}$, where $m=|V_{j}|\geq 3$. In the rest of the proof we may thus assume that $|V_{i}|\geq 3$ for $i\in[2]$.
Suppose that $V_{i}$ has at most one bipartite dominating vertex, say $v_{i,1}$ for some $i\in[2]$. Then Staller selects $v_{i,1}$ in her first move and she is able to selects her second vertex so that Dominator cannot finish the game within two moves, a contradiction. Therefore, each $V_{i}$ has at least two bipartite dominating vertices.
Suppose that $V_{i}$ for some $i\in[2]$ contains three bipartite dominating vertices, say $v_{i,1},v_{i,2}$, and $v_{i,3}$. Let $v_{j,1}$ and $v_{j,2}$ be two bipartite dominating vertices of $V_{j}$. Consider the edge $e=v_{i,3}v_{j,3}$. Then $v_{i,1},v_{j,1},v_{i,2},v_{j,2}$ are bipartite dominating vertices in $G-e$. Therefore Dominator has a strategy to win the game on $G-e$ in two moves by selecting one of the bipartite dominating vertices in each $V_{i}$, $i\in[2]$, no matter how Staller plays. Thus $\gamma_{\rm MB}^{\prime}(G-e)=2$ and hence $G$ is not $2$-$\gamma_{\rm MB}^{\prime}$-critical, a contradiction. Therefore each $V_{i}$, $i\in[2]$, has exactly two bipartite dominating vertices, say $v_{1,1},v_{2,1},v_{1,2},v_{2,2}$.
Finally suppose that there is an edge $e$ in $G$ with both of its end vertices in $V(G)\setminus\{v_{1,1},v_{2,1},v_{1,2},v_{2,2}\}$. Therefore the vertices $v_{1,1},v_{2,1},v_{1,2},v_{2,2}$ are bipartite dominating vertices in $G-e$ and hence $\gamma_{\rm MB}^{\prime}(G-e)=2$, a contradiction with $G$ being $2$-$\gamma_{\rm MB}^{\prime}$-critical. Therefore at least one end points of each edge of $G$ is a bipartite dominating vertex and thus $\deg_{G}(x)=2$ for any $x\in V(G)\setminus\{v_{1,1},v_{2,1},v_{1,2},v_{2,2}\}$. $\square$
If $n,m\geq 3$, then let $B_{n,m}$ be the bipartite graph with a bipartition $V_{1},V_{2}$, where $|V_{1}|=n$, $|V_{2}|=m$. The set $V_{1}$ contains exactly two vertices of degree $m$, the set $V_{2}$ contains exactly two vertices of degree $n$, while all the other vertices of $B_{n,m}$ are of degree $2$. This uniquely defines $B_{n,m}$. With this definition in hand the following characterization of connected bipartite 2-$\gamma_{\rm MB}^{\prime}$-critical graphs can be deduced from Lemmas 4.5 and 4.6.